OpenStudy (anonymous):

The temperature of a chemical reaction ranges between 30 degrees Celsius and 70 degrees Celsius. The temperature is at its lowest point when t = 0, and the reaction completes 1 cycle during a 10-hour period. What is a cosine function that models this reaction?

OpenStudy (anonymous):

f(t) = 20 cos 10t + 50 f(t) = 50 cos 10t + 20 f(t) = −50 cos pi over 5t + 20 f(t) = −20 cos pi over 5t + 50

OpenStudy (anonymous):

@Phebe @Liv1234 @hopeycookie @Mehek14